OpenStudy (xenesis98):

In the expression (x + y)^8, if x = 0.3 and y = 0.7, what is the numerical value of the third term? a. 0.010 b. 0.020 c. 0.058 d. 0.060 e. 0.068

MsBrains (ms-brains):

What do you think it is?

OpenStudy (xenesis98):

uh

OpenStudy (xenesis98):

a?

OpenStudy (xenesis98):

What do you think it is?

OpenStudy (reemii):

what is the formula for the third term?

OpenStudy (xenesis98):

idk

OpenStudy (reemii):

do you understand what the statement is asking you to look at?

OpenStudy (xenesis98):

no?

OpenStudy (reemii):

I think they have Newton's binomial formula in mind. Have you studied this already?

OpenStudy (xenesis98):

nope

OpenStudy (reemii):

Then you cannot answer the question… https://en.wikipedia.org/wiki/Binomial_theorem According to this formula, the 3rd term will be \[ C^8_2 (0.3)^6 (0.7)^2 = 28 (0.3)^6 (0.7)^2= 0.01000188\]

OpenStudy (xenesis98):

oh so i was right

OpenStudy (reemii):

How did you find it?

OpenStudy (xenesis98):

just a guess tbh

OpenStudy (reemii):

I encourage you to learn Newton's formula, it's a great one.. and useful.